OpenStudy (anonymous):

calculate the average rate of change for the function f(x)=-x4+4x3-2x2+x+1, from x=0 to x=1 a) 0 b) 1 c) 2 d) 7

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[ f(x)=-x^4+4x^3-2x^2+x+1\] right?

OpenStudy (anonymous):

yes.

OpenStudy (misty1212):

you need two numbers both are easy to find you need \(f(0)\) which is just \(1\) since \[f(0)=-0^4+4\times 0^3-2\times 0^2+0+1=1\]

OpenStudy (misty1212):

also you need \[f(1)=-1+4-2+1+1=3\]

OpenStudy (misty1212):

then subtract \(3-1=2\) is your answer, as always it is choice C

OpenStudy (anonymous):

oh thank you!! that was very simple and I completely understand how to do this with a different equation.